OpenStudy (anonymous):

what is the quotient of 72,000 and 90

OpenStudy (jamesj):

What's the quotient of 72 and 9? 72/9 = 8 What's the quotient of 720 and 9? 720/9 = 80 Waht's the quotient of 720 and 90? 720/90 = 8 Hopefully you can figure this out now.

OpenStudy (anonymous):

so I would do, 720 and 900?

OpenStudy (jamesj):

Given what I've shown you, what is 7,200/9 = ...?